# CSF: Fixed-outline Floorplanning Based on the Conjugate Subgradient Algorithm Assisted by Q-Learning

### v1

@@ Abstract
To perform the fixed-outline floorplanning problem efficiently, we propose to solve the original nonsmooth analytic optimization model via the conjugate subgradient algorithm (CSA), which is further accelerated by adaptively regulating the step size with the assistance of Q-learning. The objective for global floorplanning is a weighted sum of the half-perimeter wirelength, the overlapping area and the out-of-bound width, and the legalization is implemented by optimizing the weighted sum of the overlapping area and the out-of-bound width. Meanwhile, we also propose two improved variants for the legalizaiton algorithm based on constraint graphs (CGs). Experimental results demonstrate that the CSA assisted by Q-learning (CSAQ) can address both global floorplanning and legalization efficiently, and the two stages jointly contribute to competitive results on the optimization of wirelength. Meanwhile, the improved CG-based legalization methods also outperforms the original one in terms of runtime and success rate.
Index Terms:
@@
Floorplanning, fixed-ouline, nonsmooth optimization, conjugate subgradient algorithm, Q-learning
I Introduction
@@
Floorplanning is a key stage in the physical design flow of very large-scale integration circuit (VLSI) [1]. With the rapid development of integrated circuit manufacturing technology, the complexity of VLSI design has increased dramatically, and modern design usually focuses on fixed-outline floorplanning (FOFP) problems [2]. Compared to layout planning problems without fixed outline, the complexity of the FOFP problem is higher due to the additional constraint of placing a large number of modules within a given outline[3]. Additionally, the usage of intellectual property (IP) core introduces hard modules in the VLSI floorplan, making it much more challenging to optimize the fixed-outline floorplan of mixed-size modules [4].
@@
Floorplanning algorithms generally fall into two categories: floorplanning algorithms based on combinatorial optimization model (FA-COM) and floorplanning algorithms based on analytic optimization model (FA-AOM). Coding the relative positions of modules by combinatorial data structures, FA-COM formulates floorplanning problems as combinatorial optimization problems addressed by heuristics or metaheuristics [5, 6, 7, 8]. Although the combinatorial codes can be naturally converted into compact floorplans complying with the non-overlapping constraint, the combinatorial explosion contributes to their poor performance on large-scale cases. A remedy for this challenge is to implement multilevel methods, in which clustering or partitioning strategies are introduced to reduce the scale of the problem to be addressed [3]. However, the multilevel strategy typically optimizes the floorplan in a local way, which in turn makes it difficult to converge to the globally optimal floorplan of large-scale VLSI designs [9, 10].
@@
FA-AOMs address analytical floorplanning models by continuous optimization algorithms, contributing to their low complexity and fast convergence in large-scale cases [11]. A typical FA-AOM consists of two stages, the global floorplanning stage generating promising results, and the legalization stage that refines the positions of modules to eliminate constraint violations. To get a nice VLSI floorplan, the cooperative design of global floorplanning and legalization is extremely challenging due to the complicated characteristics of floorplanning models. To perform the global floorplanning efficiently, a first-order gradient-based analytic algorithm is typically incorporated to address a smoothed optimization model [12, 13], however, the smoothing approximation could introduce some distortion to the characteristics of the mathematical model. The legalization for the results of global floorplanning can be implemented via the refinement of constraint graphs [12, 13] or the second-order cone programming strategy, which cannot strike a good balance on the time complexity and the minimization of wirelength.
@@
  • The nonsmooth wirelength and overlapping metrics are directly optimized by the conjugate subgradient algorithm (CSA). Since the CSA searches the solution space by both the deterministic gradient-based search and the stochastic exploration, it is much more likely to get the optimal floorplan by flexibly regulating its search step size.
  • The CSA step size is dynamically adapted based on Q-learning, which contributes to its fast convergence to optimal floorplan.
  • Both global floorplanning and legalization are implemented by the CSA assisted by Q-learning. Consequently, the proposed fixed-outline floorplanning algorithm can strike a balance between time complexity and floorplan quality, and it demonstrates superior stability to the compared algorithms.
@@
  • Two improved variants of the CG-based legalization algorithms are proposed. They take less time to get legal floorplans with shorter wirelength.
@@
The remainder of this paper is organized as follows. Section II reviews related work. Section III introduces some preliminaries. Then, the proposed algorithm developed for fixed-outline floorplanning problems is presented in Sections IV. The numerical experiment is performed in Section V to demonstrate the competitiveness of the proposed algorithm, and finally, this paper is concluded in Section VI.
II Related Work
@@
By formulating the floorplanning problem as an analytic optimization model, Zhan et al. [14] proposed an algorithm framework consisting of the global floorplanning stage and the legalization stage, demonstrating performance superior to the Parquet-4 based on sequence pair and simulated annealing [3]. Lin and Hung [15] modelled the global floorplanning as a convex optimization problem, where modules are transformed into circles, and a push-pull mode is proposed with consideration of wirelength. Then, a second-order cone programming is deployed to legalize the floorplan. By incorporating the electrostatic field model proposed in eplace [16], a celebrated algorithm for standard cell placement in VLSI, Li et al. [12] proposed an analytic floorplanning algorithm for large-scale floorplanning cases, and horizontal constraint graphs and vertical constraint graphs were constructed to eliminate overlap of floorplanning results. Huang et al. [13] further improved the electrostatics-based analytical method for fixed-outline floorplanning, in which module rotation and sizing were introduced to get floorplanning results with shorter wirelength. Yu et al. [17] developed a floorplanning algorithm based on the feasible-seeking approach, achieving better performance than the state-of-the-art floorplanning algorithms.
@@
Since the typical wirelength and overlap metrics of analytical floorplan model are not smooth, additional smoothing approximation must be incorporated to achieve the feasibility of gradient-based optimization algorithms [18]. Popular smoothing approximation methods for the half-perimeter wirelength (HPWL) include the quadratic model, the logarithm-sum-exponential model, and the $L_{p}$-norm model, and the bell-shaped smoothing method is adopted to smooth the overlap metric [18]. In addition, Ray et al. [19] proposed a non-recursive approximation to approximate the max function for non-smooth half-perimeter wirelength models, Chan et al. [20] employed a Helmholtz smoothing technique to approximate the density function. However, the additional smoothing process not only introduces extra time complexity of the FA-AOM, but also leads to its local convergence to a solution significantly different from that of the original non-smooth model. Zhu et al. [21] proposed to address the non-smooth analytical optimization model of cell placement by the CSA, and Sun et al. [22] further developed a CSA-based floorplanning algorithm for scenarios without soft module.
@@
Recently, machine learning (ML) algorithms are introduced to generate general floorplanning algorithms. The generalization ability of well-trained ML algorithms could contribute their potential applications in diverse scenarios of floorplanning, but it is challenging to get an excellent ML-based floorplanning algorithm in the case of inadequate training data. He et al. [23] employed agents guided by the Q-learning algorithm, which alleviates the requirement of too much prior human knowledge in the search process. Mirhoseini et al. [24] proposed a RL-based method for chip floorplanning, and developed an edge-based graph convolutional neural network architecture capable of learning the transferability of chips. By merging RL with the graph convolutional network (GCN), Xu et al. [25] proposed an end-to-end learning-based floorplanning framework named as GoodFloorplan. Yang et al. [26] proposed an end-to-end reinforcement learning (RL) framework to learn a policy of floorplanning, and promising results are obtained with the assistance of edge-augmented graph attention network, position-wise multi-layer perceptron, and gated self-attention mechanism.
III Preliminaries
III-A The Non-smooth Optimization Model of Floorplanning
@@
Let $\mathbf{x}=(x_{1},x_{2},\dots,x_{n})$ and $\mathbf{\mathbf{y}}=(y_{1},y_{2},\dots,y_{n})$ be the vectors of $x$-coordinate and $y$-coordinate of all modules, respectively. The fixed-outline floorplanning (FOFP) is to place modules within a fixed outline such that all of them are mutually non-overlapping, formulated as
$\begin{array}[]{rl}\min&W(\mathbf{x},\mathbf{y})\\ s.t.&\begin{cases}D(\mathbf{x},\mathbf{y})=0\\ B(\mathbf{x},\mathbf{y})=0\end{cases}\end{array}$ (1)
where $W(\mathbf{x},\mathbf{y})$, $D(\mathbf{x},\mathbf{y})$, $B(\mathbf{x},\mathbf{y})$ are the metrics of wirelength, overlapping and bound excess, respectively. For the global floorplanning stage, we first solve the unconstraint optimization problem
$\min\,\,f_{g}(\mathbf{x},\mathbf{y})=\alpha W(\mathbf{x},\mathbf{y})+\lambda D(\mathbf{x},\mathbf{y})+\mu B(\mathbf{x},\mathbf{y})$ (2)
where $\alpha$, $\lambda$, and $\mu$ are weight coefficients. Then, the legalization is performed by addressing the prolem
$\min\,\,f_{l}(\mathbf{x},\mathbf{y})=\lambda D(\mathbf{x},\mathbf{y})+\mu B(\mathbf{x},\mathbf{y})$ (3)
In (2) and (3), we set $\mathbf{x},\mathbf{y}\in\mathbf{R}^{n}$. Then, both of them are mixed-variable optimization problems. $W(\mathbf{x},\mathbf{y})$, $D(\mathbf{x},\mathbf{y})$ and $B(\mathbf{x},\mathbf{y})$ are defined as follows.
Total Wirelength $W(\mathbf{x},\mathbf{y})$
@@
Sum of Overlapping Area $D(\mathbf{x},\mathbf{y})$
@@
Sum of Length Beyond the Fixed Outline $B(\mathbf{x},\mathbf{y})$
@@
$\begin{cases}0\leq x_{i}-\hat{w}_{i}/2,\quad x_{i}+\hat{w}_{i}/2\leq W^{*}\\ 0\leq y_{i}-\hat{h}_{i}/2,\quad y_{i}+\hat{h}_{i}/2\leq H^{*}\end{cases}$
where $W^{*}$ and $H^{*}$ are the width and the height of fixed-outline, respectively. Let
$b_{1,i}(\mathbf{x})=\max\limits_{i}(0,\hat{w}_{i}/2-x_{i}),b_{2,i}(\mathbf{x})=\max\limits_{i}(0,\hat{w}_{i}/2+x_{i}-W^{*}),$
$b_{1,i}(\mathbf{y})=\max\limits_{i}(0,\hat{h}_{i}/2-y_{i}),b_{2,i}(\mathbf{y})=\max\limits_{i}(0,\hat{h}_{i}/2+y_{i}-H^{*}).$
Accordingly, $B(\mathbf{x},\mathbf{y})$ is defined as
$B(\mathbf{x},\mathbf{y})=\sum_{i=1}^{n}(b_{1,i}(\mathbf{x})+b_{2,i}(\mathbf{x})+b_{1,i}(\mathbf{y})+b_{2,i}(\mathbf{y}))$ (8)
@@
III-B The Conjugate Subgradient Algorithm (CSA)
The pseudo code of CSA is presented in Algorithm 1. The iteration process of CSA starts with an initial solution $\mathbf{u}_{0}$ and the subgradient $\mathbf{g}_{0}$, and the conjugate direction is initialized as $\mathbf{d}_{0}=\mathbf{0}$. At generation $k$, the searching direction is confirmed according to the subgradient $\mathbf{g}_{k}$ and the conjugate direction $\mathbf{d}_{k}$, and the searching step size is determined by dividing a scaling factor $c$ by the norm of $\mathbf{d}_{k}$. When $\mathbf{u}$ locates a non-smooth point of $f$, the computation of subgradient is performed with a random scheme. Thus, the CSA is not necessarily gradient descendant. Consequently, the adaptive regulation of step size plays a critical role and has a significant influence on its convergence performance, which can be achieved by a reinforcement learning strategy.
Algorithm 1 $CSA$
@@ III-B The Conjugate Subgradient Algorithm (CSA)
III-C Q-learning
We propose to update the scaling parameter $c$ in Algorithm 1 by Q-learning, a typical value-based reinforcement learning algorithm [27]. Its goal is to find for each state-action pair $(s,a)$ a Q-value that represents its long-term expected return. The Q-values record an individual’s learning experience, stored in a Q-table and updated by
$Q^{k}(s_{j},a_{i})=(1-\alpha_{0})Q^{k-1}(s_{j},a_{i})+\alpha_{0}(R(s_{j},a_{i})+\gamma_{0}Q_{max}),$ (9)
where $Q^{k}(s_{j},a_{i})$ represents the Q-values of the state-action pair $(s_{j},a_{i})$ at steps $k$. $\alpha_{0}$, $\gamma_{0}$ are the learning rate and the discount rate, and $R(s_{j},a_{i})$ is the reward function. In this paper, $Q_{max}$ is confirmed as
$Q_{max}=m_{0}Q^{k-1}(s_{j},a_{i})+\frac{1-m_{0}}{p-1}\sum_{l\neq j}^{p}Q_{max}^{k-1}(s_{l}),$ (10)
where $m_{0}$ is a preset coefficient. $Q_{max}^{k-1}(s_{l})$ is the maximum Q-value in all actions of state $l$ in the previous iteration.
III-D The Constraint Graph
The constraint graph (CG) is a directed acyclic graph (DAG) used to represent the relative position among various modules in a layout. Modules are represented by vertexes of a CG, and arcs indicate the constraint relationships between modules. Accordingly, the horizontal constraint graph (HCG) and the vertical constraint graph (VCG) are constructed, which are composed of a series of arcs to show the positional relationships left-right and down-up between modules, respectively. Given a floorplan of modules, the HCG and the VCG can be constructed by visiting all modules. If the lower left corner coordinates of module $A$ are less than those of module $B$, then the arcs of the HCG and the VCG are constructed in the following way [12]:

### v2

@@ Abstract
The state-of-the-art researches indicate that analytic algorithms are promising in handling complex floorplanning scenarios. However, it is challenging to generate compact floorplans with excellent wirelength optimization effect due to the local convergence of gradient-based optimization algorithms designed for constructed smooth optimization models. Accordingly, we propose to construct a nonsmooth analytic floorplanning model addressed by the conjugate subgradient algorithm (CSA), which is accelerated by a population-based scheme adaptively regulating the stepsize with the assistance of Q-learning. In this way, the proposed CSA assisted by Q-learning (CSAQ) can strike a good balance on exploration and exploitation. Experimental results on the MCNC and GSRC benchmarks demonstrate that the proposed fixed-outline floorplanning algorithm based on CSAQ (CSF) not only address global floorplanning effectively, but also get legal floorplans more efficiently than the constraint graph-based legalization algorithm as well as its improved variants. It is also demonstrated that the CSF is competitive to the state-of-the-art algorithms on floorplanning scenarios only containing hard modules.
keywords:
@@
Fixed-outline floorplanning , nonsmooth analytic optimization , conjugate subgradient algorithm , Q-learning , very large-scale integration circuit
journal: Elsevier
[…]
{highlights}
The nonsmooth wirelength and overlapping metrics are directly optimized by conjugate subgradient algorithm (CSA). Since the CSA searches the solution space by both the deterministic gradient-based search and the stochastic exploration, it is much more likely to get the optimal floorplan by flexibly regulating its stepsize.
The CSA is implemented in a population-based search pattern equipped with a dynamical stepsize regulated based on Q-learning, which contributes to its fast convergence to the optimal floorplan.
Two improved variants are proposed for the constraint graph-based legalization algorithm. They take less time to get legal floorplans with shorter wirelength.
We propose a CSA-based fixed-outline floorplanning (CSF) algorithm, where both global floorplanning and legalization are implemented by the CSA assisted by Q-learning (CSAQ). It can strike a balance between time complexity and floorplan quality, and demonstrates superior performance to the compared algorithms.
1 Introduction
@@
Floorplanning is a key stage in the physical design flow of very large-scale integration circuit (VLSI) [1]. With the rapid development of integrated circuit manufacturing technology, the complexity of VLSI design has increased dramatically, and modern design usually focuses on fixed-outline floorplanning (FOFP) problems [2]. Compared to layout planning problems without fixed outline, the FOFP problem is more challenging due to the additional constraint of placing a large number of modules within a given outline [3]. Additionally, the usage of intellectual property (IP) core introduces hard modules in the VLSI floorplan, making it much more challenging to optimize the fixed-outline floorplan of mixed-size modules [4].
@@
Floorplanning algorithms generally fall into two categories: floorplanning algorithms based on combinatorial optimization model (FA-COM) and floorplanning algorithms based on analytic optimization model (FA-AOM). Coding the relative positions of modules by combinatorial data structures, FA-COM formulates floorplanning problems as combinatorial optimization problems addressed by heuristics and metaheuristics [5, 6, 7, 8]. Although the combinatorial codes can be naturally converted into compact floorplans complying with the non-overlapping constraint, the combinatorial explosion contributes to their poor performance on large-scale cases. A remedy for this challenge is to implement multilevel methods, in which clustering or partitioning strategies are introduced to reduce the scale of the problem [3]. However, the multilevel strategy typically optimizes the floorplan in a local way, which in turn makes it difficult to converge to the globally optimal floorplan of large-scale VLSI designs [9, 10].
@@
FA-AOM address analytical floorplanning models by continuous optimization algorithms, contributing to their low complexity and fast convergence in large-scale cases [11]. A typical FA-AOM consists of two stages, the global floorplanning stage generating promising results, and the legalization stage that refines the positions of modules to eliminate constraint violations. To get a nice VLSI floorplan, the cooperative design of global floorplanning and legalization is extremely challenging due to the complicated characteristics of floorplanning models. The efficient global floorplanning is typically implemented by addressing a smoothed optimization model with a first-order gradient-based analytical algorithm [12, 13], however, the smoothing approximation could introduce some distortion to the characteristics of the mathematical model, and the first-order analytical algorithm usually converges to a local optimal solution. Legalization can be implemented through the refinement of constraint graphs [12, 13] or the second-order cone programming strategy [14], which cannot strike a good balance between time complexity and minimization of wirelength.
@@
    The nonsmooth wirelength and overlapping metrics are directly optimized by CSA. Since the CSA searches the solution space by both the deterministic gradient-based search and the stochastic exploration, it is much more likely to get the optimal floorplan by flexibly regulating its step size.
  • 2.
    The CSA is implemented in a population-based search pattern equipped with a dynamical stepsize regulated based on Q-learning, which contributes to its fast convergence to the optimal floorplan.
  • 3.
@@
    Two improved variants are proposed for the CG-based legalization algorithm [12, 13]. They take less time to get legal floorplans with shorter wirelength.
  • 4.
    We propose a CSA-based fixed-outline floorplanning (CSF) algorithm, where both global floorplanning and legalization are implemented by the CSA assisted by Q-learning (CSAQ). It can strike a balance between time complexity and floorplan quality, and demonstrates superior performance to the compared algorithms.
@@
The remainder of this paper is organized as follows. Section 2 reviews related work. Section 3 introduces some preliminaries. Then, the proposed algorithm developed for fixed-outline floorplanning problems is presented in Sections 4. The numerical experiment is performed in Section 5 to demonstrate the competitiveness of the proposed algorithm, and finally, this paper is concluded in Section 6.
2 Related Work
@@
By formulating the floorplanning problem as an analytic optimization model, Zhan et al. [15] proposed an algorithm framework consisting of the global floorplanning stage and the legalization stage, demonstrating performance superior to the Parquet-4 based on sequence pair and simulated annealing [3]. Lin and Hung [14] modelled the global floorplanning as a convex optimization problem, where modules are transformed into circles, and a push-pull mode is proposed with consideration of wirelength. Then, a second-order cone programming is deployed to legalize the floorplan. By incorporating the electrostatic field model proposed in eplace [16], a celebrated algorithm for standard cell placement of VLSI, Li et al. [12] proposed an analytic floorplanning algorithm for large-scale floorplanning cases. Then, the horizontal constraint graph (HCG) and the vertical constraint graph (VCG) were constructed to eliminate overlap of floorplanning results. Huang et al. [13] further improved the electrostatics-based analytical method for fixed-outline floorplanning, in which module rotation and sizing were introduced to get floorplanning results with shorter wirelength. Yu et al. [17] developed a floorplanning algorithm based on the feasible-seeking approach, achieving better performance than the state-of-the-art floorplanning algorithms.
@@
Since the typical wirelength and overlap metrics of analytical floorplanning models are not smooth, additional smoothing approximation must be incorporated to achieve the feasibility of gradient-based optimization algorithms. Popular smoothing approximation methods for the half-perimeter wirelength (HPWL) include the quadratic model, the logarithm-sum-exponential model, and the $L_{p}$-norm model, and the bell-shaped smoothing method can be adopted to smooth the overlap metric [18]. In addition, Ray et al. [19] proposed a non-recursive approximation to simulate the max function for the non-smooth HPWL model, Chan et al. [20] employed a Helmholtz smoothing technique to approximate the density function. However, the additional smoothing process not only introduces extra time complexity of the FA-AOM, but also leads to its local convergence to a solution significantly different from that of the original non-smooth model. Zhu et al. [21] proposed to address the non-smooth analytical optimization model of cell placement by the CSA, and Sun et al. [22] further developed a CSA-based floorplanning algorithm for scenarios without soft modules.
@@
Recently, machine learning (ML) algorithms were introduced to design general floorplanning algorithms. The generalization ability of well-trained ML algorithms could contribute their potential applications in diverse scenarios of floorplanning, but it is challenging to get an excellent ML-based floorplanning algorithm in the case of inadequate training data. He et al. [23] employed agents guided by the Q-learning algorithm, which alleviates the requirement of too much prior human knowledge in the search process. Mirhoseini et al. [24] proposed a RL-based method for chip floorplanning, and developed an edge-based graph convolutional neural network architecture capable of learning the transferability of chips. By merging RL with the graph convolutional network (GCN), Xu et al. [25] proposed an end-to-end learning-based floorplanning framework named as GoodFloorplan. Yang et al. [26] proposed an end-to-end reinforcement learning (RL) framework to learn a policy of floorplanning, and promising results were obtained with the assistance of edge-augmented graph attention network, position-wise multi-layer perceptron, and gated self-attention mechanism.
3 Preliminaries
3.1 The Non-smooth Analytical Optimization Model of Floorplanning
@@
Let $\mathbf{x}=(x_{1},x_{2},\dots,x_{n})$ and $\mathbf{\mathbf{y}}=(y_{1},y_{2},\dots,y_{n})$ be the vectors of $x$-coordinate and $y$-coordinate of all modules, respectively. The FOFP aims to place modules within a fixed outline such that all of them are mutually non-overlapping, formulated as
$\begin{array}[]{rl}\min&W(\mathbf{x},\mathbf{y}),\\ s.t.&\begin{cases}D(\mathbf{x},\mathbf{y})=0,\\ B(\mathbf{x},\mathbf{y})=0,\\ \mathbf{x},\mathbf{y}\in\mathbf{R}^{n},\end{cases}\end{array}$ (1)
where $W(\mathbf{x},\mathbf{y})$, $D(\mathbf{x},\mathbf{y})$, $B(\mathbf{x},\mathbf{y})$ are the analytical metrics of wirelength, overlapping and boundary violation, respectively.
The Total Wirelength $W(\mathbf{x},\mathbf{y})$
@@
The Total Overlapping Area $D(\mathbf{x},\mathbf{y})$
@@
The Boundary Violation $B(\mathbf{x},\mathbf{y})$
@@
$\begin{cases}0\leq x_{i}-\hat{w}_{i}/2,\quad x_{i}+\hat{w}_{i}/2\leq\mathbf{W^{*}}\\ 0\leq y_{i}-\hat{h}_{i}/2,\quad y_{i}+\hat{h}_{i}/2\leq\mathbf{H^{*}}\end{cases}$
where $\mathbf{W^{*}}$ and $\mathbf{H^{*}}$ are the width and the height of fixed-outline, respectively. They can be manually assigned or generated by
$\mathbf{W^{*}}=\sqrt{(1+\gamma)\cdot A/R},\quad\mathbf{H^{*}}=\sqrt{(1+\gamma)\cdot A\cdot R},$ (6)
where $A$ is the area sum of modules, $R$ is the width-height ratio, and $\gamma$ is the ratio of whitespace.
Denote
$b_{1,i}(\mathbf{x})=\max(0,\hat{w}_{i}/2-x_{i}),b_{2,i}(\mathbf{x})=\max(0,\hat{w}_{i}/2+x_{i}-\mathbf{W^{*}}),$
$b_{1,i}(\mathbf{y})=\max(0,\hat{h}_{i}/2-y_{i}),b_{2,i}(\mathbf{y})=\max(0,\hat{h}_{i}/2+y_{i}-\mathbf{H^{*}}),$
and we define $B(\mathbf{x},\mathbf{y})$ as
$B(\mathbf{x},\mathbf{y})=\sum_{i=1}^{n}(b_{1,i}(\mathbf{x})+b_{2,i}(\mathbf{x})+b_{1,i}(\mathbf{y})+b_{2,i}(\mathbf{y})).$ (7)
For the global floorplanning stage, we first solve the unconstraint optimization problem
$\min\,\,f_{g}(\mathbf{x},\mathbf{y})=\alpha W(\mathbf{x},\mathbf{y})+\lambda D(\mathbf{x},\mathbf{y})+\mu B(\mathbf{x},\mathbf{y}),$ (8)
where $\alpha$, $\lambda$, and $\mu$ are weight coefficients. Then, the legalization is implemented by addressing the prolem
$\min\,\,f_{l}(\mathbf{x},\mathbf{y})=\lambda D(\mathbf{x},\mathbf{y})+\mu B(\mathbf{x},\mathbf{y}).$ (9)
@@
3.2 The Conjugate Subgradient Algorithm
The pseudo code of CSA is presented in Algorithm 1 [27]. The iteration process of CSA starts with an initial solution $\mathbf{u}_{0}$ and the subgradient $\mathbf{g}_{0}$, and the conjugate direction is initialized as $\mathbf{d}_{0}=\mathbf{0}$. At generation $k$, the searching direction is determined according to the subgradient $\mathbf{g}_{k}$ and the conjugate direction $\mathbf{d}_{k}$, and the searching step size is determined by dividing a scaling factor $c$ by the norm of $\mathbf{d}_{k}$. When $\mathbf{u}$ lies at a non-smooth point of $f$, the computation of subgradient is performed with a random scheme. Thus, the CSA is not necessarily gradient descendant. Consequently, the adaptive regulation of step size plays a critical role and has a significant influence on its convergence performance, which can be achieved by a reinforcement learning strategy.
Algorithm 1 CSACSA
@@ 3.2 The Conjugate Subgradient Algorithm
It was proved that the CSA presented in Algorithm 1 can converge to the neighborhood of the local optimum [27]. However, the local optimums of (8) and (9) do not guarantee the high quality of floorplanning results. To improve the global exploration without sacrificing the local exploitation, we propose to adaptively regulate the step size of CSA based on Q-learning.
3.3 Q-learning
We propose to update the scaling parameter $c$ in Algorithm 1 by Q-learning, a typical value-based reinforcement learning algorithm [28]. It aims to find for each state-action pair $(s,a)$ a Q-value that represents its long-term expected return. The Q-values record an individual’s learning experience, stored in a Q-table and updated by
$Q^{k}(s_{i},a_{j})=(1-\alpha_{0})Q^{k-1}(s_{i},a_{j})+\alpha_{0}(R(s_{i},a_{j})+\gamma_{0}Q_{max}),$ (10)
where $Q^{k}(s_{i},a_{j})$ represents the Q-values of the state-action pair $(s_{i},a_{j})$ at steps $k$. $\alpha_{0}$, $\gamma_{0}$ are the learning rate and the discount rate, respectively. $R(s_{i},a_{j})$ is the reward function. In this paper, $Q_{max}$ is confirmed as
$Q_{max}=m_{0}Q^{k-1}(s_{i},a_{j})+\frac{1-m_{0}}{p-1}\sum_{l\neq j}^{p}Q_{max}^{k-1}(s_{l}),$ (11)
where $m_{0}$ is a preset coefficient, $Q_{max}^{k-1}(s_{l})$ is the maximum Q-value across all actions of state $l$ at iteration k1k-1.
3.4 The Constraint Graph
The constraint graph (CG) is a directed acyclic graph (DAG) used to represent the relative position between modules in a layout. Modules are represented by vertexes of a CG, and arcs indicate the constraint relationships between modules. Accordingly, the horizontal constraint graph (HCG) and the vertical constraint graph (VCG) are constructed for a layout to indicate the left-right and down-up positional relationships between modules, respectively. If the lower-left corner coordinates of module $A$ are less than those of module $B$, then the arcs of the HCG and the VCG are constructed by the method proposed as below [12].